OpenStudy (anonymous):

In triangle ABC, centroid D is on median (of line) AM. Ad = x + 5 and DM = 3x - 5. Find AM

OpenStudy (anonymous):

The centroid cuts the median into the proportions 2/3 and 1/3, but the ratio of the lengths is 2:1. From your information, we have: 2(3x-5)=x+5 5x=15 x=3 so, AM= 2(3x-5)+(x+5)=2(9-5)+8=16 see: http://www.algebra.com/algebra/homework/Triangles/Triangles.faq.question.257344.html

OpenStudy (anonymous):

that's not right.. the answers provided are 12, 6, 13, and 4.

OpenStudy (anonymous):

ok. what about: AM=(5+3)+(3*3-5)=8+4=12 AM=12 honestly,... it's more on a guessing part. I'll check now & come back

OpenStudy (anonymous):

here is your problem solved & 12 is the right answer! http://answers.yahoo.com/question/index?qid=20100604120231AAfsRIT

OpenStudy (anonymous):

Great! thanks!